Find the integral surface of the pde pq=z containing the curve x-0, z-y2

1
Expert's answer
2021-01-20T12:14:34-0500

"F(z,p,q)=pq-z=0"

Trivial solution:

"z=f(x+ay)"

Let "x+ay=u"

then:

"z=f(u), \frac{\partial u}{\partial x}=1, \frac{\partial u}{\partial y}=a"

"p=\frac{\partial z}{\partial x}=\frac{dz}{du}\frac{\partial u}{\partial x}=\frac{dz}{du}"

"q=\frac{\partial z}{\partial y}=\frac{dz}{du}\frac{\partial u}{\partial y}=a\frac{dz}{du}"


Substituing values of "p" and "q" in the initial equation, we get:

"a(\frac{dz}{du})^2=z"

"\frac{dz}{du}=\sqrt{z/a}"

"du=dz/\sqrt{z/a}"

"u=2\sqrt{z/a}+b"

"x+ay=2\sqrt{z/a}+b"


If "x=0, z=y^2" , then the solution is:


"ay=2y/\sqrt{a}+b"

"y=\frac{b}{a-2/\sqrt{a}}, z=(\frac{b}{a-2/\sqrt{a}})^2"


Also:

"\frac{dz}{du}=-\sqrt{z/a}"

"x+ay=-2\sqrt{z/a}+b"

"ay=-2y/\sqrt{a}+b"

"y=\frac{b}{a+2/\sqrt{a}}, z=(\frac{b}{a+2/\sqrt{a}})^2"